What is the domain of ((x-5)/(25-(x^2)))
Answer by lwsshak3(6472)

What is the domain of ((x-5)/(25-(x^2)))
(x-5)/(25-(x^2))
When the denominator=0, the expression becomes undefined.
So, set the denominator=0 and solve for x.
25-(x^2)=0
(5+x)(5-x)=0
x≠-5 or 5
Domain: all real numbers except x=-5 and 5
(-∞,-5) U (-5,5) U (5,∞)